OpenStudy (anonymous):

Find the value of x in the interval (-π, 0) where the graph of the function fx = x+2sinx has a horizontal tangent line.

OpenStudy (anonymous):

why is nobody viewing this aahhhhhhhh

OpenStudy (anonymous):

\[f'(x)=2\cos x\]This is zero when cosx is zero. This occurs at x=-pi/2

OpenStudy (accessdenied):

I believe the derivative should be "f'(x) = 1 + 2cos x."

OpenStudy (anonymous):

ooops....made an error. should be f'(x)=1+2cosx. This gives -1/2=cosx...

OpenStudy (anonymous):

so \[x=\frac{-2\pi}{3}\]